OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL FOR BEST ANSWER Write the following inequality in slope-intercept form. -6x + 3y ≥- 45

OpenStudy (kdr1308):

Okay, so we know slope-intercept form is y = mx + b, and -6x + 3y is greater than or equal to -45 is in standard form. When you solve for x and y we end up with x = -7.5 and y = 15, so the slope is 1/2 (because 7.5 is half of 15). Now we need to plug the known values in. 15 = 1/2(7.5) + b. We don't know what b is so we need to solve it like a two-step equation to get the value of b

OpenStudy (kdr1308):

I mean greater than or = to instead or just equals sorry!

OpenStudy (mrcoolguy):

are you faze adapt

OpenStudy (kdr1308):

?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

you can first consider \[-6x+3y=-45\]

OpenStudy (anonymous):

you want \[y = mx+b\] form, where m is slope and b in intercept (i.e. x=0)